Waterways Continuing Problem 12

At the end of June the manager of the B.C. manufacturing plant was provided with the following variance analysis report:

Budget Actual Variance Favourable (F)/
Unfavourable (U)
Production in units 332,000 347,000 15,000 F
Production costs:
   Direct material $996,000 $1,017,940 $(21,940) U
   Direct labour 1,411,000 1,442,700 (31,700) U
   Variable overhead costs 166,000 172,957 (6,957) U
   Fixed overhead costs 174,300 168,620 5,680 F
Total production costs $2,747,300 $2,802,217 $(54,917) U


The manager immediately called the production supervisor, demanding an explanation for the large unfavourable variance for the quarter. The production supervisor was puzzled. He thought the cost-cutting measures they had incorporated were beginning to work. He certainly wasn’t expecting such a large discrepancy.

The standard rates the plant was using with its normal costing system are summarized below.

Volume Cost
Direct material 1.50 kg per unit $2.00 per kg
Direct labour 0.25 hour per unit $17.00 per hour
Predetermined overhead rate:
   Variable 0.25 hour per unit $2.00 per hour
   Fixed 0.25 hour per unit $2.10 per hour


Other relevant information:

1. A total of 528,000 kg of direct materials were purchased during the quarter at a cost of $2.05 per kilogram.
2. A total of 508,970 kg of direct materials were used in production to manufacture 347,000 units.
3. Payroll recorded 85,875 direct labour hours at an average cost of $16.80 per hour.


Calculate the following production variances.

Material price variance $

Unfavourable / Neither favourable nor unfavourable / Favourable

Material quantity variance $

Unfavourable / Favourable / Neither favourable nor unfavourable

Labour price variance $

Unfavourable / Neither favourable nor unfavourable / Favourable

Labour efficiency variance $

Favourable / Neither favourable nor unfavourable / Unfavourable

Variable overhead variance $

Favourable / Neither favourable nor unfavourable / Unfavourable
Answer #1

Material price variance = (Standard Price – Actual price)*Actual quantity purchased

= (2-2.05)*528000

= $26,400 Unfavorable

Material Quantity Variance = (Standard Quantity – Actual Quantity Used)*Standard Price

= (347000*1.5 – 508,970)*2

= $23,060 Favorable

Labor price variance = (Standard Rate – Actual Rate)*Actual Hours

= (17-16.80)*85875

= $17,175 Favorable

Labor Efficiency variance = (Standard Hours – Actual Hours)*Standard Rate

= (347000*0.25 – 85875)*17

= $14,875 Favorable

Variable overhead variance = Standard Cost – Actual Cost

= 347000*0.25*2 – 172957

= $543 Favorable